The desired daily output for an assembly line is 540 units. This assembly line will operate 729 minutes per day. The following table contains information on this product's task times and precedence relationships: TASK TASK TIME (SECONDS) IMMEDIATE PREDECESSOR A 30 — B 28 A C 18 B D 69 B E 34 A F 36 E G 42 C, D H 24 F, G b. What is the workstation cycle time? Cycle time seconds c. Balance this line using the largest number of following tasks. Use the longest task time as a secondary criterion. (Leave no cells blank - be certain to enter "0" wherever required.) Workstation Task Idle Time I II III IV d. What is the efficiency of your line balance? (Round your answer to 2 decimal places.) Efficiency %
Answer #1

Task

Task time- secs

Predecessor

A

30

-

B

28

A

C

18

B

D

69

B

E

34

A

F

36

E

G

42

C, D

H

24

F, G

Answer b: desired output = 540 units per 729 mins= 729*60/540= 81 secs per units

Hence the workstation cycle time required is 81 secs per unit

Answer c: Line balancing, required cycle time is 81 secs per workstation. Sum of task times in a workstation cannot be more than cycle time of 81 secs.

Workstation

Tasks

Idle time= Cycle time- sum of task time

Reason

1

A, B, C

81-(30+28+18) = 5

Start with A also include B- largest following tasks, there is time space for C also

2

D

12

D has more following tasks than E

3

E, F

11

Assign E first and then F

4

G, H

15

Only tasks remaining

Answer d: Balanced delay = no. of workstation*cycle time – total task time/ no. of workstation* cycle time =(4*81- 281)/(4*81)= 0.1327 ie 13.27%

Efficiency= 1-balanced delay= 86.73%